OpenStudy (anonymous):

How do I write an equation in slope intercept form of a line that passes through (-2, 10), (2,2), & (4,-2)?

OpenStudy (anonymous):

Slope intercept form is (y - a) = m(x - b), where m is the slope and (a, b) is a point that is given. You are given three points, so choose any of the two that you want. Let's choose the first two. The slope is the change in y over the change in x, so let's use (-2, 10) and (2, 2). The change in y is 10 - 2 = 8. The change in x is -2 - 2 = -4. So the slope is 8/(-4) = -2. Plugging into our equation, (y- a) = m (x - b) becomes (y - a) = -2(x - b). For a and b, you can use any of your points. So correct answers are: y - 10 = -2(x +2) y - 2 = -2 (x - 2) y + 2 = -2 (x - 4)

OpenStudy (anonymous):

But are trying to say y -y1 = m (x -x1)? Thats the one I use

OpenStudy (anonymous):

It's the same equation, yes. Just different variables used. But yes, use that general form!

OpenStudy (anonymous):

Ok! Thnx

OpenStudy (anonymous):

You're very welcome!